OpenStudy (anonymous):

PLEASE HELP<3 as you know i give medalsss Solve y = 3x – 12 by the substitution method. 2x – 3y = -13

OpenStudy (bahrom7893):

GOO IGBA!

OpenStudy (anonymous):

You guys are so supportive ;p

OpenStudy (bahrom7893):

lol we take turns haha

OpenStudy (anonymous):

arent they lol :p

OpenStudy (anonymous):

I love it, it cracks me up

OpenStudy (lgbasallote):

2x - 3(2x - 12) = - 13 2x - 6x + 36 = - 13 -4x = -49 x = 49/4 is that right bahrom? i'm having doubts because of the awkwardness of the improper fraction =)))

OpenStudy (bahrom7893):

y = 3x – 12 2x – 3y = -13 2x - 9x + 36 = -13 -7x = -49 x = 7 y = 21-12 = 9

OpenStudy (bahrom7893):

right?

OpenStudy (lgbasallote):

ahhh.. i knew my answer was awkward =))) i mistyped and miscalculated haha...that's right bahrom

OpenStudy (anonymous):

lol yes bathroom thats right and ya ilikegoodsmlls i get alot of laughsfrom them :p

OpenStudy (lgbasallote):

HAHAHAHAHA =))))))))))))))

OpenStudy (anonymous):

wellll then y is ur name bathroom? wtf would u like me to call u

OpenStudy (anonymous):

my b

OpenStudy (anonymous):

lol i did a min ago

OpenStudy (anonymous):

lol i did a min ago

OpenStudy (anonymous):

lol bathroom :p too funny my BAHROM

OpenStudy (anonymous):

ayee ok look i swear it said bathroom